Fermi National Laboratory


Accelerator Update for September 19, 2005 - September 21, 2005

Monday September 19
The day shift began with Operations monitoring store 4392 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Operations began preparing the Recycler to stash antiprotons at 10:58 AM.

At 11:19 AM, Preaccelerator experts accessed the H- Source to replace an extractor tube. They completed the job about twenty minutes later.

The Recycler stash ended at 11:42 AM.

Pbar resumed stacking and then at 11:59 AM experts began a reverse proton study.

MCenter resumed taking beam at 12:05 PM.

At 12:16 PM, a Booster RF power supply (BRF11) tripped off. An observer saw smoke and the Fire Department responded. After securing the hazard, the Fermi Fire Captain turned the area over to Operations. A LCW hose leak had shorted out a transformer. Operations called in RF experts to replace the transformer.

At 12:50 PM, Operations reported that BRF18 was off and a Booster expert was investigating. BRF12 also tripped off, but it reset okay.

Operations reported at 2:40 PM that BRF11 was back on line.

At 2:59 PM, the Booster notcher tripped off for the first of two times in thirty-eight minutes. It reset both times.

Pbar experts completed their study and Pbar resumed stacking at 4:37 PM.

At 4:50 PM, Operations lost their network connection to the NuMI BPMs. NuMI beam was halted. Operators investigated and found water dripping from a clogged air conditioner drain into the rack that housed the network hub for NuMI controls. An operator called in an expert.

At 5:06 PM, MI RF stations 11, 14, and 16, and then later 10 tripped off. An RF expert made some adjustments. They all reset okay.

At 5:34 PM, the Booster notcher tripped off. It reset okay.

Operations reported at 6:07 PM that NuMI was back to its normal intensity.

At 6:20 PM, MIRF15 tripped off for the first of many times in about two hours. Operators left the station off after its last trip.

At 8:10 PM, the Booster notcher tripped off. Operators had it reset eight minutes later.

The Recycler began preparing to stash at 9:33 PM. The stash ended at 10:41 PM.

Pbar resumed stacking and MCenter resumed taking beam 10:51 PM.

Around 11:40 PM, the Booster notcher tripped off. It reset okay.

Tuesday September 20
The midnight shift began with Operations monitoring store 4392 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 3:18 AM, the Booster notcher tripped off. It reset okay.

TeV experts began an end-of-store study around 5 AM.

Operations terminated store 4392 at 7:14 AM.

Tevatron (TeV) experts began a study at 7:49 AM.

Operations put the TeV through a dry squeeze at 12:01 PM.

Operations put the TeV into shot setup around 1 PM.

Preaccelerator experts began switching from the H- Source to the I- Source at 2:35 PM. They completed the switch 11 minutes later.

Operations established Recycler-only store 4395 at 3:08 PM with an initial luminosity of 90.4E30. Pbar resumed stacking and MCenter resumed taking beam.

The Recycler began preparing to stash antiprotons around 4:07 PM. They completed the stash and Pbar resumed stacking at 5:17 PM. MCenter resumed taking beam.

The Recycler began stashing at 9:06 PM and ended at 9:19 PM. Pbar and MCenter resumed normal operations at 9:35 PM.

Wednesday September 21
The midnight shift began with Operations monitoring store 4395 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:30 AM, the TeV suffered a sector D1 quench. Operators investigated and found that a recent reboot set minimum magnet helium flow values lower that they should have been.

At 1:48 AM, a Linac quadrupole bulk power supply tripped off for the first of many times over the next four hours due to a high temperature indication. It reset each time.

Operators began the TeV turn on at 6:11 AM.

Operations put the TeV into a dry squeeze at 6:48 AM.

The Recycler began preparing for stashing at 7:52 AM.

The Plans for Wednesday
The plans for today are to stack, stash, and store. MCenter will be off during the day shift for scheduled work in the Meson tunnels.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/21/2005   email Fermilab